### v1

# Heralded Emission Detection in Quantum Dot Solids under Twin-Photon Excitation

Abstract
Harnessing quantum correlations between photons is an emerging direction in optical spectroscopy, but experimental progress has thus far been limited to fluorescence-detected studies of molecular systems at room temperature. Here, we demonstrate heralded emission detection (HED) under continuous-wave (CW) entangled-photon excitation of near-infrared (NIR)–emitting colloidal III–V quantum dots (QDs) at low temperatures. We use superconducting nanowire single-photon detectors (SNSPDs) and large-area NIR avalanche photodiodes (APDs) to maximize the time resolution ($\sim$70 ps) and emission count rates ($\sim$2000 cps). To our knowledge, these results represent the first demonstration of HED in the solid state and establish NIR QDs as a bright, tunable model system with longer-lived coherences and integrability with optical cavities as an additional experimental control knob.
1 Introduction
Optical spectroscopy continues to illuminate foundational processes in chemical, biomolecular, and solid-state systems across timescales. Thus far, canonical optical spectroscopy uses classical light from pulsed or continuous-wave (CW) lasers producing coherent wavepackets or wavetrains. The conventional semiclassical description for the light-matter interaction, treating light as a classical perturbation of a quantum system, is therefore typically sufficient [1, 2]. Electronic or vibrational coherences generated through pulse/sample interaction can be interfered with additional time-delayed optical pulses, thereby bestowing non-linear spectroscopy with exquisite control via the bandwidth, shape, and delay between pulses [3]. These control knobs can extract the energy-time correlations of a sample response without requiring any quantization of field excitations, i.e., without invoking photons.
A relatively new paradigm in optical spectroscopy uses quantum light i.e., photons with defined correlations in time, energy, or polarization [4, 5]. This is distinctly different from classical spectroscopy, where coherent laser light is characterized by a Poissonian number distribution of otherwise uncorrelated photons [6, 7]. Quantum spectroscopy can therefore use what classical spectroscopy cannot; the inherent correlations between photons in bespoke quantum states. Recent theoretical efforts have explored non-classical correlations between photons in spectroscopic experiments e.g., using energy-time entangled pairs. Intriguing theoretical predictions [4] beyond overcoming classical noise limits have been made [8]. For instance, access to classically forbidden excitation pathways [9], control over photo-chemical conversion rates through vibrational wavepacket preparation [10, 11], or remote-controlled spectroscopy via twin-pair detection of entangled states have been proposed [12].
Only a few experimental studies on entangled-photon light-matter interaction follow the theoretical lead. As such, the optical response of materials and molecules under quantum light excitation, e.g., single-photon Fock states or energy-time entangled photon pairs from spontaneous parametric downconversion (SPDC) has gradually come into focus [13, 14], but is still limited by the significant experimental complexity of quantum light spectroscopy. In this context, the still-limited brightness of entangled photon sources used for sample excitation is a particular challenge [15]. This is different from most spectroscopic experiments, in which sample stability rather than the achievable excitation power is the limiting factor. Probing sample coherences after entangled-photon interaction with the sample using a ’probe’ pulse (or second photon) is therefore not easily achieved due to the low probability of two photons interacting with the same emitter. Instead, either the entangled photon phase delay (in transmission) or spontaneous emission as proxy for entangled photon absorption is measured.
Two common entangled two-photon (ETP) spectroscopies include fluorescence-detected two-photon absorption [16, 17, 18] and heralded emission detection (HED) [19, 20, 21, 22, 23]. Two-photon absorption using ETP has been reported to exhibit stronger absorption compared to classical two-photon absorption, owing to the intrinsic time correlation of photons in pairs [16]. This effect has potential applications in fluorescence microscopy [17], although it remains somewhat debated [24]. HED leverages the intrinsic time correlation of photon pairs from quantum light sources such as SPDC. One photon excites the emissive sample, while its twin is directly detected to “herald” the excitation event. Correlating the herald with the following photoluminescence photon reconstructs the emission dynamics, using the twin-pair correlations. This is fundamentally different from the laser-pulse correlations employed in conventional lifetime measurements using time-correlated single-photon counting (TCSPC). HED lifetime measurements therefore work with pulsed and CW lasers. The wavelengths of the entangled photons can be finely tuned by adjusting the temperature of the nonlinear crystal in the SPDC process, allowing for wavelength-dependent excitation studies. Because SPDC produces photon statistics in each arm that follow a thermal distribution, HED can be employed to study emission lifetimes under thermal light-like conditions. This makes it especially valuable for exploring systems such as photosynthetic complexes or photovoltaic materials under illumination with realistic underlying photon statistics. Another key difference is that TCSPC requires very low detection probability per pulse (1–5%) to avoid pile-up, which limits usable counts [25]. HED, not tied to an excitation clock, avoids pile-up and can in principle operate at higher fluxes. Despite its conceptual simplicity, only a few studies have applied HED, all of which have measured solution-phase molecular systems such as Indocyanine Green (ICG) [19], Rhodamine 6G (R6G) [20], and IR-140 [23]. A noteworthy feature of HED is its ability to herald single-photon Fock states to study light-driven processes in the ultimate limit of excitation intensity, e.g., in Light-Harvesting 2 (LH2) complexes [21, 22].
To advance quantum spectroscopy, both the scope of accessible methods and samples must be expanded. HED’s current main limitation, shared with all other entangled-pair spectroscopy, is the low photon flux and coincidence rates from entangled-photon sources. HED is therefore currently not competitive for routine lifetime measurements. However, it can provide a foundation for extending quantum-light spectroscopy beyond simple intensity correlations, opening the possibility to resolve additional energy- and polarization-correlations. Such capabilities would enable more sophisticated approaches for probing quantum systems, with the broader aim of accessing entanglement signatures and unraveling excitonic dynamics in quantum materials [26]. In light of these requirements, the expansion of entangled-pair spectroscopy from solution-phase organic molecules to solid-state materials is a logical next step. First, the entangled light-matter interaction and emission can be photonically enhanced to increase both the overall brightness and excitation rate per emitter. The latter is needed, for example, to validate theoretical predictions of inter-pair delay-dependent two-photon excitation i.e., electronic pathway selectivity [27]. To advance quantum spectroscopy, both the scope of accessible methods and samples must be expanded. HED’s current main limitation, shared with all other entangled-pair spectroscopy, is the low photon flux and coincidence rates from entangled-photon sources. HED is therefore currently not competitive for routine lifetime measurements. However, it can provide a foundation for extending quantum-light spectroscopy beyond simple intensity correlations, opening the possibility to resolve additional energy- and polarization-correlations. Such capabilities would enable more sophisticated approaches for probing quantum systems, with the broader aim of accessing entanglement signatures and unraveling excitonic dynamics in quantum materials [26]. In light of these requirements, the expansion of entangled-pair spectroscopy from solution-phase organic molecules to solid-state materials is a logical next step. First, the entangled light-matter interaction and emission can be photonically enhanced to increase both the overall brightness and excitation rate per emitter. The latter is needed, for example, to validate theoretical predictions of inter-pair delay-dependent two-photon excitation i.e., electronic pathway selectivity [27].
Here, we demonstrate HED in solid-state materials at low temperatures, specifically thin-films of near-infrared (NIR) emitting colloidal InAs/ZnSe quantum dots (QDs). QDs offer a rich platform for exploring entangled-pair light–matter interactions, thanks to their tunable electronic fine structure, phonon coupling, and electron–hole radiative and non-radiative rates [28]. To our knowledge, our results represent the first demonstration of entangled-photon heralded emission detection in a solid-state material, distinct from prior two-photon excitation approaches using entangled photons [29]. QD’s processability and proven integration with nanophotonic structures may help overcome brightness limitations through enhancement of the absorption and emission [30], and serve as a model system for entangled-photon non-linear spectroscopy.
2 Result
We developed an improved HED apparatus capable of operating over variable solid-state sample temperatures, as shown in Fig. 1a. A detailed description of the setup components is provided in the Supporting Information. Briefly, entangled photon pairs are generated via SPDC by pumping a periodically poled potassium titanyl phosphate (ppKTP) crystal with a 405 nm CW laser under the type-II phase-matching condition. The generated pairs exhibit a bandwidth of $\sim$2 nm with spectral overlap consistent with degenerate SPDC (Fig. 1b). A Hong–Ou–Mandel (HOM) visibility of 86% confirms the high indistinguishability of the photon pairs, in agreement with previous reports (Fig. S1) [31]. The two photons are separated by a fiber-coupled polarization beam splitter (PBS), yielding a heralding arm and a sample arm. The heralding photons are directly detected by an SNSPD, while the sample-arm photons are sent through a 50:50 beamsplitter into a cryostat with a 0.9 numerical aperature (NA) cryo-objective to excite the sample. The resulting emission is collected by the same objective, filtered from the excitation photons, and coupled via a fiber collimator into a second SNSPD. The use of SNSPDs is a notable feature of our apparatus due to their high photo-detection efficiency in the NIR ($\sim$80%), minimal dark counts (<<10 Hz), and low dead time ($\leq$35 ns). Their additionally low timing jitter (31 ps and 50 ps) provides an overall narrow ($\sim$72 ps) instrument response function (IRF) for coincidence counts without sample (Fig. 1c). The IRF width is mainly determined by the convolution of the timing jitter of the SPDC source, SNSPDs, and the single-photon counting module (19 ps).
We selected colloidal InAs/ZnSe core–shell QDs as the model system for HED. These QDs are promising NIR emitters[32] and are also widely applied in photonic devices such as photoconductors[33], phototransistors[34], and photodiodes[35, 36]. Importantly, colloidal InAs QDs can be directly deposited onto photonic platforms[37], offering a straightforward route toward integration with quantum photonic structures for optical communication and emerging quantum photonic applications.
Figure 1: Experimental setup, source characteristics, and properties of InAs/ZnSe nanocrystals. (a) Experimental setup for heralded emission detection and preparation of InAs/ZnSe quantum dot (QD) films. (b) Normalized spectra of the entangled photons from Type-II SPDC used in the heralding and sample arms. (c) The impulse response function (IRF), obtained by replacing the sample with a cover slip coated with a thin gold film and collecting the reflected photons. (d) Ensemble absorption and emission spectra of InAs/ZnSe QDs dispersed in toluene at room temperature. The shaded region marks the twin-pair excitation wavelength in HED. For clarity, the excitation line is drawn with exaggerated width and does not represent the actual narrow bandwidth. Inset: representative transmission electron microscopy (TEM) images of InAs/ZnSe QDs. Scale bar: 30 nm.
Our synthesized type-I heterostructured InAs/ZnSe core–shell nanocrystals display high (approximately 58%) ensemble photoluminescence quantum yield (PLQY) in solution, consistent with previously reported values for this material system [38]. This PLQY is substantially higher than of organic dyes in the NIR [39], rendering NIR-QDs a natural choice for HED with high signal strengths. Room temperature QDs ensemble absorption and emission spectra reveal a low prominence excitonic shoulder around 899 nm based on Elliott model fitting (Fig. S3) and a redshifted, broad emission around 984 nm, shown in Fig. 1d. We attribute the absence of narrower features to inhomogeneous broadening due to size- and shape-inhomogeneity, which is common even for the currently best InAs QDs [32, 40]. The transmission electron microscopy (TEM) image, as shown in the inset of Fig. 1d, confirms the tetrahedral QD shape [38] and a mean particle size of 8.2 $\pm$ 2.2 nm (Fig. S3). This size is only a fraction of the InAs bulk Bohr radius of around 35 nm [41], rendering our QDs well within the strong confinement regime, where size dispersion has a high impact on spectral broadening[32].
For emissive thin-film preparation, we employed drop-casting onto gold-coated coverslips (Fig. 1a) rather than conventional spin coating. Spin-coated films typically yield thin layers ($\sim$30–100 nm) [42, 43], which contain only a limited number of QDs within the optical focal volume (Rayleigh length × focus spot size), thereby reducing the probability of photon absorption. Consequently, the detected HED signals can be weak even for materials with high PLQY. One strategy to increase film thickness is to mix QDs with polymers such as polystyrene or poly(methyl methacrylate). However, achieving a thickness comparable to the Rayleigh length requires a high polymer concentration, which increases the polymer-to-QD ratio within the focal volume, reduces film brightness, and may ntroduce non-negligible quenching effects (see Supporting Information). In principle, higher QD loadings could offset this issue, but in practice such concentrations are difficult to achieve. Another potential approach is layer-by-layer deposition, in which oleic acid–capped QDs are sequentially coated, the surface wetting properties are modified via solid-state ligand exchange, and additional layers are deposited. While this method can produce multilayered QD films [44], it often comes at the expense of reduced PLQY and requires further optimization [43]. By contrast, drop-casting naturally produces thicker and denser films, thereby improving absorption efficiency and yielding stronger HED signals. The drawback, however, is that drop-cast films typically exhibit non-uniform coverage, as shown in Fig. S3. Representative SEM images, also in Fig. S3, further highlight their inhomogeneous morphology. This inhomogeneity makes it challenging to systematically study the relationship between film brightness and thickness.
Figure 2: Data analyses in heralded emission detection. (a) In heralded emission detection (HED), photon pairs are split into a heralding arm (blue) and a sample arm (orange). Pairs can be classified as coincidence pairs, which are time-correlated and produce bunching in the $g^{(2)}(\tau)$, or non-coincidence pairs, which contribute uncorrelated background. The sample emission therefore possesses a similar statistics with emission photon from both twins (red) and accidental pairs (light red). (b) Intensity traces of the heralding arm and the emission signal during a 20h measurement at 4.7 K, demonstrating stable down-conversion over a ten-hour period. (c) Normalized cross-correlation $g^{(2)}(\tau)$ between the heralding arm and the emission at 4.7 K. The orange-shaded region corresponds to contributions from correlated heralding–emitted photon pairs (entangled two-photon down-conversion), while the purple region represents uncorrelated background with $g^{(2)}(\tau)=1$. (d) Schematic for the $g^{(2)}$ correlation analysis of time-tagged HED data. Delaying one sequence relative to the other only shifts the bunching peak without changing its shape. (e) Schematic of the effect of adopting a ’start–stop’ alternative from time-correlated single-photon counting (TCSPC) to HED, introducing distortions.
Fig. 2 illustrates the concept, stability of emission, and different approaches for analyzing HED data. Three photon “arms” are considered in the measurement: the heralding arm, the sample arm used to excite the sample, and the emission arm, which collects photons emitted by the sample following excitation. SPDC sources are not ideal [45]; in addition to generating true coincidence photon pairs, non-coincidence pairs are also produced, as shown in Fig. 2a. Consequently, the emitted sample photons originate from both non-coincidence and coincidence pairs. The former produces accidental heralding-emission coincidences. The cample emission count rates remained stable over the course of the experiment ($\sim$15–18 hours), confirming the sample’s stability under low excitation fluences (see Fig. 2b for 4.7 K and Fig. S4 for all temperatures). At 4.7 K, the heralding arm consistently recorded an average count rate of $\sim 3.2\times 10^{6}$ counts per second (cps). Assuming the sample arm exhibits a similar rate, and accounting for losses through optical elements, the number of excitation photons reaching the sample is estimated to be $1.12\times 10^{6}$, corresponding to a photon power of $\sim$275 fW. Emitted photons were collected at a rate of $\sim$89 cps (see Supporting Information for all temperatures and details).
We now analyze correlations between heralding and sample photons, extending approaches previously applied to molecular dyes at room temperature [19, 21, 20, 23, 22]. Leveraging the short 70 ps IRF of our SNSPDs, we resolve the normalized cross-correlation $g^{(2)}(\tau)$ between emission and heralding photons (Fig. 2c). As expected, a distinct bunching peak >1>1 at $\tau=0$ indicates preservation of the bunched statistics of twin pairs in the heralded emission. The background with $g^{(2)}(\tau)=1$ results from accidental coincidence events. The decay of $g^{(2)}(\tau)>1$ with $\tau$ measures the effective delay between heralded photon absorption and emission, i.e. the spontaneous emission lifetime [19]. We confirm the near-equivalency of HED lifetimes of our QDs and TCSPC, as shown in the Supporting Information. Small deviations are most likely due to differences in the excitation wavelength or small deviations in setup alignments between different excitation sources. Interestingly, we identify a consistent dip of $g^{(2)}<1$ for $\tau<0$ (Fig. 2c, inset). Such anti-correlations would be consistent with time reversal symmetry breaking, but likely have a more trivial origin. We assign the origin of this dip to a non-obvious artifact from detector deadtime. Coincidences with $\tau<0$ are from emitted photons leading heralding photons. A leading emitted photon has a larger than random probability of being led by a second, even earlier heralding photon. The detection of this second, earliest, heralding photon turns the heralding arm SNSPD dark deadtime (<<20 ns). The number of following emission/heralding coincidences within the deadtime is thus reduced, producing the dip in $g^{(2)}(\tau<0)$. This artifact may be misinterpreted or may go unnoticed, especially for longer deadtime detectors. While the above correlation analysis is established, we like to point out that translation of the “start-stop” paradigm used in pulsed-laser TCSPC would require correction in post-processing and is more prone to distortion from pathlength differences. We clarify this in Fig. 2d-e.
The start–stop algorithm selects a photon on one channel as “start” and a photon on the other channel as “stop,” storing their time differences in a histogram. By contrast, the correlation algorithm accumulates all time differences between detection events on the two channels into a histogram, treating every click as both “start” and “stop,” and calculating both positive and negative delays. Using a start–stop scheme in HED can introduce distortions, since the overall number of pairs along the lifetime not only depends on the heralding ratio and entanglement time, but also the waiting time distribution of the accidental pairs. This can produce spurious accidental coincidences that are difficult to separate from the lifetime decay. Timing misalignment between emitted photon and their heralding photon can further deform the decay dynamics (Fig. 2d). The $g^{(2)}$ correlation approach avoids these complications, as it does not impose fixed start/stop assignments on the photons. Any time delay between the heralding and emission arms merely shifts the correlation peak without distorting its decay dynamics (Fig. 2e). Waiting time distributions between accidental counts are naturally factored out in the normalized $g^{(2)}$.
One shortcoming of many SNSPDs is their single-mode fiber access [46], causing substantial insertion losses. If high count rates rather than minimal timing jitter are desirable in HED, large area APDs are preferred. Fig. 3a–c showcases HED results acquired across different temperatures using both an SNSPD (Fig. 3a) and a large-area APD (Fig. 3b-c). Although the employed APD exhibited high dark counts ($\sim$1848 cps), they contribute only to accidental coincidences and do not affect the shape of the bunching peak. We consistently observed total emission count rates of $\sim$4000 cps when accounting for accidental dark counts. Sample 1, with a quantum dot size of 8.2 nm, and Sample 2, with a size of 6.3 nm, showed similar absolute brightness despite their differing lifetimes. These detected count rates led to an increased external quantum efficiency (EQE) for the film i.e., the ratio of detected emission photons to heralded sample arm photons (Fig. 3f).
Figure 3: Temperature-dependent HED with SNSPD and APD for sample emission detection (a) Normalized cross-correlation $g^{(2)}(\tau)$ at various temperatures tagging the sample emission with an SNSPD. The bunching peak reflects time correlations governed by the photoluminescence lifetime of the InAs/ZnSe quantum dots (QDs). (b) Normalized $g^{(2)}(\tau)$ of a second QD sample measured with an avalanche photodiode (APD). (c) Data from (b) on a logarithmic scale. (d) Extracted lifetimes as a function of temperature, along with bright–dark model fits. The inset illustrates the model, where G, B, and D denote the ground, bright, and dark state, respectively. (e) Brightness and bunching peak area extracted from normalized $g^{(2)}(\tau)$ data as a function of temperature. The bunching area is fitted with a Bose–Einstein model. (f) External quantum efficiency (EQE) of the twin photon downconversion by the sample at different temperatures, following the trend in brightness.
To demonstrate the HED analysis with these count rates, we used it to quantify temperature-dependent lifetimes. Similar to sample 1, all decay traces consistently display at least two components, independent of temperature. With increasing temperature, the fastest decay component, prominent at low temperatures, gradually diminishes. This trend is not caused by the gold-coated substrate, as control measurements on nanocrystals deposited on plain microscope slides also exhibited the same fastest decay component, as demonstrated for sample 1 (Fig. S5). The initial fast decay component in all $g^{(2)}$ traces is consistent with the rapid decays previously reported in InSb [47, 48] and InAs/CdSe [49] nanocrystals. Such behavior is commonly attributed either to nonradiative recombination [47, 48] or to relaxation from the bright to the dark exciton state [49]. The slower decay components are interpreted in terms of thermally activated dark-to-bright state population [50, 51], corresponding to the average lifetimes of the bright and dark exciton states. The whole lifetimes were extracted using a biexponential model for $T\leq 52.0$ K, where one component captures the initial fast decay and the other represents the average lifetimes of the bright and dark states. At higher temperatures, where the fast component is absent, a single-exponential fit was applied to maintain consistency. This average lifetime can be interpreted as arising from two excitonic states in thermal equilibrium, separated by an energy splitting $\Delta E$. The dark state exhibits a longer lifetime, $\tau_{D}$, whereas the bright state has a shorter lifetime, $\tau_{B}$, in agreement with the bright–dark exciton model.
$\tau^{-1}_{avg}=\left(\frac{e^{\Delta E/k_{B}T}}{1+e^{\Delta E/k_{B}T}}\right)\tau^{-1}_{D}+\left(\frac{1}{1+e^{\Delta E/k_{B}T}}\right)\tau^{-1}_{B}$ (1)
The fitting results (Fig. 3d) yield $\tau_{D}\approx 228.3$ ns, $\tau_{B}\approx 16.6$ ns, and $\Delta E\approx 6.7$ meV, values comparable to those reported for InP/ZnSe QDs [52]. The discrepancies between the data and the fit at 6.3 K, likely arise from either insufficient separation of the fastest decay component during lifetime extraction or from the need for a more comprehensive model that explicitly includes the fast component along with relaxation between bright and dark states or other contributing processes.
Fig. 3e further display the relationship between the sample brightness and the total area under the $g^{(2)}(\tau)$ bunching peak in Fig. 3b, which both diminish with temperature. We point out that the decrease in brightness qualitatively matches a Bose-Einstein model for phonon population with a characteristic energy of $\approx 65~meV$, consistent with phonon energies reported for III–V QDs [52, 53], possibly suggesting phonon-mediated turn-on of non-radiative decay. This effect was less pronounced in Sample 1 (Fig. S9).
3 Discussion
Advancing quantum spectroscopy will require both improved apparatus capabilities and broader material platforms. This study contributes by identifying colloidal InAs/ZnSe QDs as bright NIR model systems and by detailing heralding analysis, detector tradeoffs, and unexpected artifact sources to guide future improvements. The main limitation in quantum spectroscopy generally is the low coincidence rate after sample interaction [19, 21]. Pinpointing losses to external quantum efficiency across setups [46]and materials is a logical next step. Given the tunability and integrability of QDs as photonic building blocks [54], their targeted optimization could deliver disproportionate gains.
Table 1 compares key parameters and achieved coincidence rates of all HED studies known to the authors, all but Eshun et al. reported low - at most around 120 -coincidence counts per second. One limitation of organic molecules is their generally decreasing luminescent quantum yield (QY) with decreasing emission energy. Few dyes in the near- and shortwave infrared have appreciable QY [39], naturally limiting the coincidence rate for molecules excitable with established sources typically operating in the NIR. Eshun et al. is a notable exception, producing 488–550 nm photons in the sample arm of non-degenerate SPDC pairs, enabling excitation of Rhodamine 6G with an appreciable quantum yield (95%). While this emission efficiency is certainly contributing to the suggested coincidence rate, we point out that there is currently no agreed on best practice for calculating these rates.
Our NIR QDs, used in conjunction with our setup, represent the brightest NIR samples applied to HED to date. We were unable to further improve brightness through systematic thin-film preparation, including multilayer spin-coating and the use of polymeric additives. Importantly, simply maximizing the thin-film PLQY is insufficient to predict the highest HED response, as brightness is also limited by film thickness and morphology. Thicker films can lead to reabsorption losses, while inhomogeneous films may scatter or trap emission, reducing out-coupling efficiency [55]. Thus, optimization of both absorption and emission into Gaussian-like modes, along with minimization of reabsorption in thicker films, is essential to achieve maximum HED brightness.
It is instructive to identify the limiting factors in the downconversion efficiency. Beyond the intrinsic loss from the objective’s collection efficiency, an important factor is the sample configuration—specifically, the number of QDs within the focal volume. Unlike organic dyes, QDs are physically larger, which lowers the probability of absorbing all incident photons, even though they typically exhibit higher PLQY. In addition, when using SNSPDs, the collection efficiency is constrained by the single-mode fibers [46]. While these fibers provide localized information from the film, they also introduce substantial count losses because the emission beam is imaged onto the fiber core, whereas the QD emission across the thin film is spatially inhomogeneous.
In our configuration—using an SNSPD in the heralding arm and a large-area NIR APD in the emission arm—we achieved high coincidence count rates of $\sim$230 coincidence counts per second (ccps) at the low temperatures. This setup maximizes detection efficiency for both herald and emitted photons while avoiding fiber-coupling insertion and losses from the incoherent sample emission into non-Gaussian modes. The trade-off, however, is the APD’s larger timing jitter, which broadens the instrument response function and reduces the precision of $g^{(2)}(\tau)$. These results highlight the importance of detector choice in entangled-photon spectroscopy and suggest that advances in quantum light sources and detectors will further benefit studies of quantum light–matter interactions [15].
Overall, this study establishes colloidal QDs as a solid-state platform for exploring entangled photon light–matter interactions and for developing twin-photon detectors and entangled-photon downconversion devices. While our data do not resolve any effects of prolonged sample coherences on entangled-photon absorption, future phase-sensitive studies—such as Franson-type interferometry [56]—could directly probe entanglement preservation and benefit from the reported brightness herein.
Table 1: Summary of HED work to date
Work Sample Lifetime (ns) Pump EPP (ccps) Excitation (cps) Emission (cps) Coincidences (ccps)
Harper et al. [19] ICG 0.62 CW $1\times 10^{5}$ 860 6.5
Li et al. [21] LH2 1.20 Pulsed $2.7\times 10^{5}$ $9.13\times 10^{5}$ 121
Eshun et al. [20]* R6G 3.99 CW $1.6\times 10^{6}$ 23000
Gäbler et al. [23] IR-140 0.93 CW
Alvarez-Mendoza et al. [22] LH2 1.14 CW $2\times 10^{5}$
This work** InAs/ZnSe QDs 40–250 CW $6\times 10^{5} 1.1$$1\times 10^{6}$ 20–4200 4–270
Note: EPP refers to the coincidence counts per second (ccps) for entangled photon pairs. Coincidences represent counts detected between the heralding and emission arms. For comparison with the literature, we report the maximum recorded values of these counts.
*The definitions of EPP and coincidences in some previous works may differ.
**Approximations or ranges are provided for each parameter to reflect variations under different experimental conditions.
4 Conclusions
We have demonstrated heralded emission detection (HED) under entangled-photon excitation in InAs/ZnSe quantum dot films at cryogenic temperatures, establishing a solid-state platform for this spectroscopy. By benchmarking against TCSPC, we show that HED reliably recovers excitonic lifetimes and bright–dark state dynamics, validating its extension beyond molecular systems. Our results highlight both opportunities and limits. QDs provide brightness, tunability, and device compatibility that make them promising for quantum-light spectroscopy. While this work produced higher coincidence rates than all but one previous work [20], and the highest in the NIR spectral range, the observed coincidence rates are still modest. The demonstrated integration of QDs with optical cavities and waveguides presents a new control knob in quantum light spectroscopy, positioning QDs as a versatile model system.
5 Methods
5.1 Synthesis of InAs/ZnSe QDs
InAs/ZnSe QDs were synthesized following previously reported methods [57, 58, 59]. Both InAs nanoclusters and InAs QDs were prepared using the same precursor.
5.1.1 Materials
[…]
5.1.2 Synthesis of InAs Clusters
[…]
5.1.3 Synthesis of InAs QDs
[…]
5.1.4 Synthesis of ZnSe shell
[…]
5.2 InAs/ZnSe QDs characterization
The measurements of UV-vis absorption were performed on a Jasco V-570 UV-vis-NIR spectrophotometer. Photoluminescence spectra and photoluminescence quantum yield was measured by a Hamamatsu Absolute PL quantum yield spectrometer C11347-11 and Edinburgh FLS920 fluorescence spectrometer. Transmission electron microscopy (TEM) images of the InAs/ZnSe QDs were acquired using a Tecnai 12 TEM. Scanning electron microscopy (SEM) images of the QDs film were obtained with a Zeiss XB 550 FIB-SEM.
Acknowledgments
We thank Tim Rambo and Ryan Wilson for helpful discussions on instrumentation. This work was financially supported by the College of Chemistry at the University of California, Berkeley and the US Department of Energy (DE-AC02-05CH11231).
Supplemental document
The data supporting the findings of this study are available within the article and its Supplementary Information.
Supplementary materials
S1 The detailed configuration of the setup
S1.1 Heraled Emission Detection (HED)
The entangled photon pair source (customized QES 2.4, Qubitekk) consists of a 405 nm continuous-wave (CW) laser that pumps a periodically poled potassium titanyl phosphate (ppKTP) crystal housed in a temperature-controlled oven, allowing for tuning of the photon wavelength and degree of degeneracy. The indistinguishability of photon pairs has been shown in Fig. S1. Entangled photon pairs are generated via spontaneous parametric down-conversion (SPDC), after which the residual pump light is filtered out. The entangled photons are then coupled into optical fiber. A polarization-maintaining, fiber-coupled polarizing beam splitter (PBS) with FC/PC connectors (Qubitekk) is connected to the source output to separate the signal and idler photons. One output of the PBS is routed via an FC/PC sleeve (ADAFC2, Thorlabs) to a single-mode fiber patch cable (780HP, Diamond), which is terminated with an E2000 connector and connected to a superconducting nanowire single-photon detector (SNSPD, Quantum Opus), referred to as the heralding arm.
The other output of the PBS is coupled into a protected-silver-coated reflective fiber collimator (RC02FC-P01, Thorlabs), producing a collimated beam with a $1/e^{2}$ diameter of approximately 1.8 mm. The beam is launched into free space, passes through an 810 nm bandpass filter (FBH810-10, Thorlabs) and protected-silver mirrors (PF10-03-P01, Thorlabs), and is elevated by a periscope assembly (RS99, Thorlabs) equipped with the same mirror type. It then passes through a 50:50 beam splitter (BS014, Thorlabs) and is directed onto an x–y galvanometric mirror system (GVS012, Thorlabs) controlled by a data acquisition (DAQ) platform (T7, LabJack). The reflected beam is subsequently relayed into a high-numerical-aperture (NA) objective (EC Epiplan-Neofluar 100×/0.9 BD DIC M27, Zeiss) housed in a closed-cycle cryostat (Montana Instruments) using a pair of relay lenses with focal lengths of 150 mm (AC254-150-AB-ML, Thorlabs) and 200 mm (AC254-200-AB-ML, Thorlabs). This configuration, together with the beam divergence, expands the spot to fill the objective’s back aperture, which has a conjugate plane size of approximately 3 mm.
Photoluminescence emission from the sample is collected by the same objective and travels back through the same relay lens pair and galvo system. The emission is separated from the excitation path by the same 50:50 BS and filtered by an 850 nm long-pass filter (FELH0850, Thorlabs) to remove residual excitation photons. The filtered emission is then collected by another fiber coupler (PAF2P-A10B FiberPort, Thorlabs) and routed via single-mode fiber patch (1060HI, Diamond) to a second SNSPD, referred to as the emission arm.
Different fibers and detectors are used for the heralding and emission arms to match their respective wavelength ranges. The Heralding arm uses 780HP fiber, which operates effectively in the 780–970 nm range, and is connected to an SNSPD with a timing jitter of approximately 31 ps, less than 1 dark count per second, and a dead time of 20 ns. The Sample arm uses HI1060 fiber, paired with an SNSPD optimized for the emission wavelength of InAs/ZnSe nanocrystals. This detector has a timing jitter of approximately 50 ps, fewer than 10 dark counts per second, and a dead time of 18 ns.
Detection events from both SNSPDs are time-stamped with 8 ps resolution using a single-photon counting module (Time Tagger Ultra, Performance Edition, Swabian Instruments). For second-order correlation measurements, the system accumulates the time differences between photon detection events on the two channels into a histogram. Each detection event is considered as both a “start” and a “stop,” allowing the calculation of both positive and negative time delays. A time bin width of 1 ns is chosen to balance the trade-off between signal-to-noise ratio (SNR) and timing resolution.
For the APD-based version of the HED, the short-pass dichroic mirror (FF825-SDI01, Semrock) replaces the PBS output path, and the signal is detected using an APD (id120, IDQ).
It should be noted that the timing jitters here are specified as full width at half maximum (FWHM).
Figure S1: Hong–Ou–Mandel (HOM) dip of the entangled photon pairs, indicating their indistinguishability, with a visibility of approximately 86%.
Table S1: Summary of detectors used in this work.
Detector Deadtime (ns) dark count (cps) Timing jitter (ps) detection efficiency
SNSPD (heralding) <20<20@780 nm <10<10 31 83%@780nm
SNSPD (emission) $\leq 35$@1315nm <10<10 50 83%@1064 nm
APD 400 $\sim$1848 <<400 10%@1000 nm
S1.2 Time-Correlated Single Photon Counting (TCSPC)
The setup for TCSPC measurements closely follows that used for HED, with the primary difference being the excitation source: a 470 nm pulsed laser (LDH-P-C-470, PicoQuant). The laser repetition rate is reduced to 2.5 MHz using an external controller (PDL 820, PicoQuant) to meet the desired excitation conditions. Along the excitation path, a 500 nm short-pass filter (FESH0500, Thorlabs) is combined with a set of neutral density filters (NE20B, NE20A, and NE10A-A; Thorlabs), reducing the excitation fluence to 6.37 nJ/cm², which cause the exciton number per pulse $\langle N\rangle\ll 1$. All other optical components remain unchanged from the HES setup.
For TCSPC measurements, a start-stop mode is used on the same Time Tagger in place of a second-order correlation algorithm. In this mode, detection events on two channels are recorded as start and stop signals, and the time differences between them are compiled into a histogram. For consistency with HED data, the histogram bin width is set to 1 ns.